I have 20 surveys and my average is 84% how many 100% (perfect surveys) do i need to get my average above 91%
Answer by Boreal(15235)

average = total/number.
Average is 0.84, number is 20, so total is 16.8
Let x = number of perfect surveys needed.
{16.8 +x}/(20+x) = 0.91
multiply by (20+x) both sides
16.8 +x = 18.2 +0.91 x
-1.4 =-0.09 x
need between 15 and 16 surveys, so round up to 16 surveys.
Check
36 will be the number of surveys.
The total will be 16.8+12=32.8
32.8/36=0.9111 (repeat)
Try 35 surveys (15 more)
31.8/35 =0.9086.
You need 16 more surveys.